OpenStudy (anonymous):

How can you represent the terms of a sequence explicitly? How can you represent them recursively?

OpenStudy (anonymous):

@ganeshie8 @hartnn @SolomonZelman please help

OpenStudy (anonymous):

@hba please help

OpenStudy (hba):

Do you know about Fibonacci sequence?

OpenStudy (anonymous):

No

OpenStudy (hba):

Do you have a particular sequence?

OpenStudy (anonymous):

no not really, just a simple explanation

OpenStudy (anonymous):

just not getting it at all

OpenStudy (hba):

0,1,1,2,3,5,8,13,21........

OpenStudy (hba):

How would you represent this sequence?

OpenStudy (anonymous):

idk 0+1=1 1+1=2 2+1=3 3+2=5

OpenStudy (hba):

See there you are :D ,So what are you doing here? :) How would you represent this :D

OpenStudy (anonymous):

3+5=8 5+8=13 13+8=21

OpenStudy (anonymous):

So thats what the question is asking? What about the "recursive" part?

OpenStudy (hba):

f(n)=f(n-1)+f(n-2)

OpenStudy (anonymous):

oh kay so now im lost again lol what does that have to do with the sequence?

OpenStudy (hba):

0[(f(n-2)],1[(f(n-1)],1[f(n)] 1=1+0

OpenStudy (anonymous):

oh and thanks for helping! lol idk how lost id be without you!

OpenStudy (hba):

HAHA lol :p

OpenStudy (anonymous):

omg you really lost me there XD 0[(f(n-2)],1[(f(n-1)],1[f(n)] WHAT?!

OpenStudy (hba):

Lol i think you are just tired and bored :p

OpenStudy (anonymous):

no im literally confused XD

OpenStudy (anonymous):

i don't understand it, i get the first part but recursive idk

OpenStudy (hba):

You need some rest believe me :p you have done way too much of math :D Allow me to explain you what recursion is first :D

OpenStudy (anonymous):

okay! its really early here lol please explain

OpenStudy (hba):

Lol it's too late for me then :D Anyways,let's get back to the thing.Have you ever done programming btw? Do you know what factorial is?

OpenStudy (anonymous):

yupppp

OpenStudy (hba):

yup for both?

OpenStudy (anonymous):

factorial yes, programming eh no lol

OpenStudy (hba):

Lol okay... :p n!=n*(n-1)*(n-2)....*1 let’s think about the factorial operation. The factorial operation is a perfect candidate for recursion because it is a problem that can easily be broken up into similar smaller problems! 5! = 5 * 4 * 3 * 2 * 1 So if we want to calculate the factorial of 5, we can think of it as multiplying 5 by the factorial of 4: 5! = 5 * 4! Similarly, to calculate the factorial of 4, we multiply 4 by the factorial of 3: 4! = 4 * 3! And so on... So if we try to calculate this we can: f(n)=n*f(n-1) if i input 3 for instance factorial(3)=3*factorial(2) factorial(3)= 3* (2*factorial(1)) factorial(3)=3*2

OpenStudy (hba):

Now you understand lol? :P

OpenStudy (anonymous):

Ah a lot better lol thanks!

OpenStudy (hba):

I know it ain't a easy concept Jelia. :D Anyways if you get rid of everything do message me i am bored :p Your welcome :)

OpenStudy (anonymous):

oh i've got more questions for ya lol give me 5 min

OpenStudy (hba):

0_0 why you do dis to me? :p

OpenStudy (anonymous):

Hey! you said ur bored! lol sooooo ill keep ya busy :) im a good friend XD

OpenStudy (hba):

Haha lol no keep me engaged by talking to me :p or doing something crazy for me :D

OpenStudy (anonymous):

i've got 5 questions for you after that ill be fun Jelia! lol

OpenStudy (hba):

fun Jelia? HAHAHAHAH :p Seriously? :D Okay stop being so funny :D

OpenStudy (anonymous):

generate the first five terms in the sequence using the explicit formula. y(base)n=-5n-5